OpenStudy (buttermequeasy):

Choose the best description of how to calculate f[g(−4)] if f(x) = 2x + 1 and g(x) = x2 − 10. Calculate f(−4) and g(−4), then multiply them together. Calculate f(−4) and g(−4), then write the results as an ordered pair. Calculate g(−4), plug the result into f(x) and simplify. Calculate f(−4) and g(−4), then subtract them.

OpenStudy (anonymous):

It's the third one. f(g(x)) means that we evaluate the x in the f function as the g(x). That is, suppose f = x^2 and g = x^3. Then, f(g(x)) = (x^3)^2